Assume that adults have IQ scores that are normally distributed with a mean of 96.8 and a standard deviation of 24.4. Find the probability that a randomly selected adult has an IQ greater than 122.9.
Answer #1

Solution :

Given ,

mean = = 96.8

standard deviation = = 24.4

P(x >122.9 ) = 1 - P(x< 122.9)

= 1 - P[(x -) / < (122.9-96.8) /24.4 ]

= 1 - P(z <1.07 )

Using z table

= 1 - 0.8577

= 0.1423

probability= 0.1423